OpenStudy (anonymous):

3^(2p) + 3^(3q) + 3^(5r) = 3^(7s) Find the minimum value of p+q+r+s where p,q,r,s are all positive integer.

OpenStudy (mayankdevnani):

9^p+27^q+3^5r=3^7s

OpenStudy (mayankdevnani):

ok

OpenStudy (mayankdevnani):

got it @sauravshakya

OpenStudy (anonymous):

What was that?

OpenStudy (mayankdevnani):

we just square and cube the value

OpenStudy (anonymous):

The question is to find minimum value of p+q+r+s

OpenStudy (mayankdevnani):

then common 3 as a base

OpenStudy (mayankdevnani):

and then add powers

OpenStudy (mayankdevnani):

then you solve it

OpenStudy (anonymous):

I have never seen that rule.

OpenStudy (mayankdevnani):

ok so go through wolframalpha

OpenStudy (anonymous):

I dont think that will also help here.

OpenStudy (mayankdevnani):

go through

OpenStudy (anonymous):

There is nothing.

OpenStudy (anonymous):

I am seeing it.

OpenStudy (anonymous):

So, what is the answer?

OpenStudy (anonymous):

Give a min, but at tell u the answer can't be find the above way.

OpenStudy (anonymous):

What I figured out till now is s must be an odd number....

OpenStudy (anonymous):

@experimentX PLZ see this

OpenStudy (experimentx):

this doesn't look nice ... saying that find the minimum value of p+q+r+s <--- this should be close to first instances.

OpenStudy (anonymous):

Cant we find that......

OpenStudy (experimentx):

exponential diophantine equations are awful!! you should probably tag mukushla

OpenStudy (anonymous):

This question is not of mine........ Someone posted this today and said a 8th grade challenging question.

OpenStudy (anonymous):

But he is not online.

OpenStudy (anonymous):

@mukushla

OpenStudy (experimentx):

you don't do diophantine equations on 8th grade.

OpenStudy (anonymous):

Even I dont know that.

OpenStudy (anonymous):

This surely cant be a 8th grade question.

OpenStudy (mayankdevnani):

@mathslover i think he do this

OpenStudy (mayankdevnani):

very dhinchach question

OpenStudy (anonymous):

Now, what I am trying to PROVE is 3^(2p) + 3^(3q) + 3^(5r) = 3^(7s) is never true for any value of p,q,r,s

OpenStudy (experimentx):

anyway this was the output I got from mathematica \[ \left\{\left\{p\to \frac{8}{5}+2 i,q\to \frac{223}{10}+\frac{53 i}{5},r\to -\frac{71}{10}-\frac{49 i}{10},\\ s\to \frac{180 i \pi +\text{Log}\left[3^{-\frac{71}{2}-\frac{49 i}{2}}+3^{\frac{16}{5}+4 i}+3^{\frac{669}{10}+\frac{159 i}{5}}\right]}{7 \text{Log}[3]}\right\}\right\} \] I trolled enough ... not time to go strolling!!

OpenStudy (anonymous):

I dont know it even has a solution.

OpenStudy (experimentx):

*now .. seeya later!!

OpenStudy (anonymous):

ok...... BYE thanx for trying.

OpenStudy (experimentx):

call it hardly a try .. hehehe

OpenStudy (shubhamsrg):

how about this : i dont know if this'll work : ( 3^(2p) + 3^(3q) + 3^(5r) )/3 >= 3^( (2p + 3q + 5r)/3 ) or 3^(7s-1) >= 3^( (2p + 3q + 5r)/3 ) ??

OpenStudy (shubhamsrg):

@ujjwal p,q,r,s are only positive integers..

OpenStudy (anonymous):

@shubhamsrg ( 3^(2p) + 3^(3q) + 3^(5r) )/3 >= 3^( (2p + 3q + 5r)/3 ) HOW???

OpenStudy (shubhamsrg):

from AM>=GM

OpenStudy (anonymous):

Oh >= got it.

OpenStudy (anonymous):

Yeah nice one.

OpenStudy (shubhamsrg):

and continuing that, we have 21s >= 2p + 3q + 5r + 3 but still no idea if that'll help..

OpenStudy (anonymous):

And I got: